OpenStudy (anonymous):

Can you solve the inequality of -50 < 5n + -5 < 30

OpenStudy (ash2326):

We have the inequality \[-50<5n+-5<30\] it's actually \[-50<5n-5<30\] I'll guide you, would you follow?

OpenStudy (anonymous):

Sure

OpenStudy (ash2326):

Add 5 to all the sides

OpenStudy (anonymous):

Can you show me how to do it, then I'll figure it out on the next one.

OpenStudy (ash2326):

sure ! \[-50<5n-5<30\] Add 5 to all the sides \[-50+5<5n-5+5<30+5\] so we get \[-45<5n+0<35\] or \[-45<5n<35\] Now we want the range in terms of n, so divide all sides by 5 \[\frac{-45}{5}<\frac{5n}{5}<\frac{35}{5}\] we get \[-9<n<7\] this is our answer

OpenStudy (anonymous):

Okay, Thank you. I believe I understand now(:

OpenStudy (ash2326):

Good:D